OpenStudy (anonymous):

separable by homogenuous equation dy/dx=(x+2y)/(3y-2x) ijust could solve until this xdv/dx=1+4V-3v^2/3v-2... can anyone comtinue to solve this???

OpenStudy (abb0t):

Um, you could use substitution for this method.

OpenStudy (abb0t):

\[x\frac{ dv }{ dx }+v = \frac{ x+2xv }{ -2x+3xv } \Rightarrow xv'+v=\frac{ x+2xv }{ -2x+3xv }\]

OpenStudy (abb0t):

\[\frac{ dv }{ dx }= \frac{ -3v^2+4v+1 }{ x(3v-2) }\]

OpenStudy (abb0t):

Now is seperable.